OpenStudy (anonymous):

y varies inversely with x, and x = 7 when y = 35. Find the inverse variation equation and use it to find the value of x when y = 50.

OpenStudy (anonymous):

yes they are; 4.9 5 0.2 250

OpenStudy (nurali):

x = 7 when y = 35. Find the inverse variation equation and use it to find the value of x when y = 50. y = (1/k)(x) [1/k because it's inversely proportional] 7 = (1/k)(35) 1/k = 35/7 = 5 y = (5)(50) y = 250

OpenStudy (anonymous):

thanks a bunch

OpenStudy (nurali):

My Pleasure.